Consider a 2nd order differential equation
d
2
dx2
u(x) = f(x), 0 ≤ x ≤ a . (1)
Find the solutions of the above equation with the boundary conditions u(0) = 0 and u(a) = 0 using the
Green’s function technique.
1
Expert's answer
2020-02-24T10:01:37-0500

As per the given question,

"\dfrac{d^2u(x)}{dx^2}+u(x)=f(x)" ",0\leq x\leq a"

let "u(x)=A (x)\cos kx+B (x)\sin kx" ----(i)

now, taking the differenciation twice with respect to x equation (i) with

"-kA'\sin kx+kB' \cos kx=f(x)"

"-k^2(A'\cos kx+B' \sin kx)=u''(x)"

So, "A'(x)=\dfrac{f(x)\sin kx}{k}" and "B'(x)=\dfrac{-f(x)\cos kx}{k}"

Now, We can write this equation as per the below

"u(x)=\dfrac{\sin kx}{k}\int^x_af(y)\sin ky dy-\dfrac{\cos kx}{k}\int^x_af(y)\cos ky dy"

Now u(0)=0

"\int^0_af(y)\sin ky dy=0"

So a=0

similarly,

"\int^x_af(y)\cos ky dy=0"

so, u(a)=0

Hence, it is satisfying the condition.